OpenStudy (anonymous):

f(x)=x2-3x-7 find f(-3) a- -7 b- 8 c- 11 d- 22

OpenStudy (anonymous):

This is so easy

OpenStudy (anonymous):

f(x) = x^2 -3x - 7 f(3) = 3^2 - 3(3) - 7 = ....

OpenStudy (anonymous):

f(3) simply means everywhere you see x, replace it with 3

OpenStudy (anonymous):

then evaluate

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

it's f(-3), not f(3)

OpenStudy (anonymous):

f(-3) = -3^2 -3(-3) - 7 = 9 + 9 - 7 = 11

OpenStudy (anonymous):

The answer is c

OpenStudy (anonymous):

oh thanks man, i am puting another question if you can take a look to it

OpenStudy (anonymous):

But try to solve it yourself first, then get to the part you are stuck on, then only post that part.